OpenStudy (anonymous):

An isosceles triangle has area of 125 ft^2. If the base is 14 ft, what is the length of each leg? Round your answer to the nearest tenth. 19.2 ft 15.8 ft 514.9 ft 22.7 ft

OpenStudy (cwrw238):



OpenStudy (cwrw238):

area = 7*h 125 = 7 * h h = 125 / 7 = 17.87 length of each leg = sqrt(7^2 + 17.87^2) = 19.2 ft

OpenStudy (anonymous):

Thank you so much!

OpenStudy (cwrw238):

you r welcome